# Lowers bounds for Waldschmidt constants and Demailly’s Conjecture for general and very general points

### v1

@@ Abstract.
We prove Demailly’s Conjecture concerning the lower bound for the Waldschmidt constant in terms of the initial degree of the second symbolic powers for any set of generic points or very general points in ${\mathbb{P}}^{N}$. We also discuss the Harbourne-Huneke Containment and the aforementioned Demailly’s Conjecture for general points and show the results for sufficiently many general points and general points in projective spaces with low dimensions.
@@ 1. Introduction
Acknowledgments. The authors thank Huy Tài Hà for his comments and suggestions. The second author is partially supported by the NAFOSTED (Vietnam) under the grant number 101.04-2023.07.
@@ 2. Preliminaries
$m^{\times s}=\underbrace{(m,m,\dots,m)}_{s\text{ times }}.$
@@
If $\widehat{\alpha}(s):=\widehat{\alpha}(I(1^{\times s}))$ is the Waldschmidt constant the defining ideal of $s$ generic points in ${\mathbb{P}}^{N}$, then the followings are true
@@
Our proof of Demailly’s Conjecture when $m=2$ relies on the lower bound for the Waldschmidt constant of a set of generic points and the upper bound of the Castelnouvo-Mumford regularity and the initial degree of the second symbolic power of the defining ideal of the set of generic points. To study the former, we utilize the Cremona reduction process, see for instance, [Dum12, Dum15, BN22], and gluing results from the Waldschmidt decomposition technique given in [DSS24]. For the latter, the celebrated Alexander-Hirschowitz Theorem provides bounds for the regularity value and the initial degree of the second symbolic power. We recall these results here.
Definition 2.6.
The Cremona transformation is the following birational map
$\Phi:{\mathbb{P}}^{N}_{\mathbbm{k}}\to{\mathbb{P}}^{N}_{\mathbbm{k}},\text{ defined by }\Phi(x_{0}:\dots:x_{N})\mapsto(x_{0}^{-1}:\dots:x_{N}^{-1}).$
We recall some useful results from Cremona reductions.
@@
We also have the following useful result whose proof using the Waldschmidt decomposition.
@@
Lastly, we recall the celebrated Alexander-Hirschowitz Theorem proven in a series of work [Ale88, AH92b, AH92a, AH95], and explain how it gives the regularity value and the initial degree of the second symbolic powers.
@@ 3. Demailly’s Conjecture when $m=2$ for Very General Points
The result for very general points follows from specialization the result for generic points, see [FMX18]. Hence, we will assume that $I$ is the defining ideal of $s$ generic points. Moreover, we will prove the (stronger) results for $N=3,4$ in Section 5, hence, we will assume here that $N\geqslant 5$. By Lemma 3.1, it suffices to show that
@@
In fact, the above inequality is equivalent to $(N+\ell)(N+\ell-1)\geqslant(N+1)\ell(\ell-1)$, which is equivalent to
$\ell^{2}-3\ell-(N-1)\leqslant 0,$
hence, is equivalent to $\dfrac{3-\sqrt{4N+5}}{2}\leqslant\ell\leqslant\dfrac{3+\sqrt{4N+5}}{2}$. Thus, the claim follows since we assume $\ell\geqslant 2>\frac{3-\sqrt{4N+5}}{2}$.
@@
($\dfrac{{N+\ell\choose N}}{N+1}\geqslant\dfrac{{N-1+\ell\choose N-1}}{N}+\dfrac{{N-1+\ell-1\choose N-1}}{N}+1.$
The inequality is equivalent to
$\dfrac{(N+\ell-2)!}{\ell!(N+1)!}(\ell^{2}-3\ell-N+1)\geqslant 1.$
@@
Thus, for $\ell>\dfrac{3+\sqrt{4N+5}}{2}$, if (${N+\ell\choose N}\leqslant(N+1)s$, since
($\dfrac{{N+\ell\choose N}}{N+1}\geqslant\dfrac{{N-1+\ell\choose N-1}}{N}+\dfrac{{N-1+\ell-1\choose N-1}}{N}+1,$
we have that
@@ 5. Ideal Containment and Demailly’s Conjecture for General Points in Low Dimension
In this section, we shall show Demailly’s Conjecture when $m=2$ for any set of general points in ${\mathbb{P}}^{3},{\mathbb{P}}^{4}$ and ${\mathbb{P}}^{5}$, except four numbers of points in ${\mathbb{P}}^{5}$ where the result is still not known. Note that the results in this section are used for the base cases in the proof of Theorem 3.2. In the following, as the conjecture was shown to be true in several cases: small number of general points in [NT19] and sufficiently many general points in [BGHN22a], we will only consider remaining cases.
@@
where $I$ is the defining ideal of $s$ generic points with (${4+\ell\choose 4}<5s\leqslant{4+\ell+1\choose 4}$. We again consider case by case as follows.
@@
Applying Lemma 2.7 for three times with largest possible multiplicities proves the claim. Indeed, each time the reductor $k=3d-\sum_{i=1}^{5}m_{i}$ affects the coefficients of $m$ in $m_{i}$ and $d$ significantly. In the following table we capture the coefficient of $m$ in $k$, and change of coefficients of $m$ in $m_{i}$ and $d$. The change in the constant term does not affect so we omit them in the table.
@@
From the last row, since, the coefficient of $m$ in the degree $d$ is less than that of the fifth point, we conclude that $I\left((20m)^{\times 9},(30m)^{\times 1}\right)_{36m-1}=0,\forall m$. Now, by applying the Lemma 2.8 one time, we get $I((20m)^{\times 15})_{36m-1}=0$, for all $m$, therefore, $\widehat{\alpha}(I)\geqslant\widehat{\alpha}(15)\geqslant\dfrac{36}{20}=\dfrac{5+4}{5}$.
@@
We claim that $I\left((20m)^{\times 1},(30m)^{\times 11}\right)_{48m-1}=0,\forall m.$ Similarly, we apply Lemma 2.7 for two times with largest possible multiplicities proves the claim. The following table we capture the coefficient of $m$ in $k$, and change of coefficients of $m$ in $m_{i}$ and $d$. Again, the change in the constant term does not affect so we omit thoses in the table.
@@
From the last row, since, the coefficient of $m$ in the degree $d$ is less than that of the first point, we conclude that $I\left((20m)^{\times 1},(30m)^{\times 11}\right)_{48m-1}=0,\forall m.$ Now, by applying the Lemma 2.8 eleven times, we get $I((20m)^{\times 67})_{48m-1}=0$, for all $m$, therefore, $\widehat{\alpha}(I)\geqslant\widehat{\alpha}(67)\geqslant\dfrac{48}{20}=\dfrac{5+7}{5}$.
@@
Combined results in [BGHN22a, DSS24, NT19], Demailly’s Conjecture when $m=2$ now holds for any set of very general points, any set of at least $2^{N}$ general points, and any set of points in ${\mathbb{P}}^{3},{\mathbb{P}}^{4}$, and ${\mathbb{P}}^{5}$, except for $10\leqslant s\leqslant 13$ in ${\mathbb{P}}^{5}$.

### v2

# Lower bounds for Waldschmidt constants and Demailly’s Conjecture for general and very general points

@@ Abstract.
We prove Demailly’s Conjecture concerning the lower bound for the Waldschmidt constant in terms of the initial degree of the second symbolic powers for any set of very general points in ${\mathbb{P}}^{N}$. We also discuss the Harbourne-Huneke Containment and the aforementioned Demailly’s Conjecture for general points and show the results for sufficiently many general points and general points in projective spaces with low dimensions.
@@ 1. Introduction
Acknowledgments. The authors thank Huy Tài Hà for his comments and suggestions. The second author (TTN) is partially supported by the NAFOSTED (Vietnam) under the grant number 101.04-2023.07. Part of this work was completed when the second author was visiting the Vietnam Institute of Advanced Study in Mathematics (VIASM). He thanks VIASM for the hospitality and support. The authors also thank the anonymous referees for their detailed comments, which helped improve the presentation of the manuscript.
@@ 2. Preliminaries
$I\left(m^{\times s}\right)=I(\underbrace{m,m,\dots,m}_{s\text{ times }}).$
@@
If $\widehat{\alpha}(s):=\widehat{\alpha}(I(1^{\times s}))$ is the Waldschmidt constant of the defining ideal of $s$ generic points in ${\mathbb{P}}^{N}$, then the followings are true
@@
Our proof of Demailly’s Conjecture when $m=2$ relies on the lower bound for the Waldschmidt constant of a set of generic points and the upper bound of the Castelnouvo-Mumford regularity and the initial degree of the second symbolic power of the defining ideal of the set of generic points. To study the former, we utilize the Cremona reduction process, see for instance, [Dum12, Dum15, BN22], and gluing results from the Waldschmidt decomposition technique given in [DSS24]. For the latter, the celebrated Alexander-Hirschowitz Theorem provides bounds for the regularity value and the initial degree of the second symbolic power. We first recall some useful results that can be proved by using Cremona transformations.
@@
We also have the following useful result whose proof uses the Waldschmidt decompositions.
@@
Lastly, we recall the celebrated Alexander-Hirschowitz Theorem, which was proved in a series of work [Ale88, AH92b, AH92a, AH95], and explain how it gives the regularity value and the initial degree of the second symbolic powers.
@@ 3. Demailly’s Conjecture when $m=2$ for Very General Points
The result for very general points follows from specializing the result for generic points, see [FMX18]. Hence, we will assume that $I$ is the defining ideal of $s$ generic points. Moreover, we will prove the (stronger) results for $N=3,4$ in Section 5, hence, we will assume here that $N\geqslant 5$. By Lemma 3.1, it suffices to show that
@@
In fact, the above inequality is equivalent to
$(N+\ell)(N+\ell-1)\geqslant(N+1)\ell(\ell-1)\iff\ell^{2}-3\ell-(N-1)\leqslant 0\iff\dfrac{3-\sqrt{4N+5}}{2}\leqslant\ell\leqslant\dfrac{3+\sqrt{4N+5}}{2}$
Thus, the claim follows since we assume $\ell\geqslant 2>\frac{3-\sqrt{4N+5}}{2}$.
@@
($\dfrac{{N+\ell\choose N}}{N+1}\geqslant\dfrac{{N-1+\ell\choose N-1}}{N}+\dfrac{{N-1+\ell-1\choose N-1}}{N}+1\iff\dfrac{(N+\ell-2)!}{\ell!(N+1)!}(\ell^{2}-3\ell-N+1)\geqslant 1.$
@@
Thus, for $\ell>\dfrac{3+\sqrt{4N+5}}{2}$, if (${N+\ell\choose N}\leqslant(N+1)s$, since ($\dfrac{{N+\ell\choose N}}{N+1}\geqslant\dfrac{{N-1+\ell\choose N-1}}{N}+\dfrac{{N-1+\ell-1\choose N-1}}{N}+1,$ we have that
@@ 5. Ideal Containment and Demailly’s Conjecture for General Points in Low Dimension
In this section, we shall show Demailly’s Conjecture when $m=2$ for any set of general points in ${\mathbb{P}}^{3},{\mathbb{P}}^{4}$ and ${\mathbb{P}}^{5}$, except four numbers of points in ${\mathbb{P}}^{5}$ where the result is still not known. Note that the results in this section are used for the base cases in the proof of Theorem 3.2. Since Conjecture 1.2 is known to be true for small number of general points [NT19] and sufficiently many general points [BGHN22a], we will only consider the remaining cases as discussed below.
@@
where $I$ is the defining ideal of $s$ generic points with (${4+\ell\choose 4}<5s\leqslant{4+\ell+1\choose 4}$.
We again consider case by case as follows.
@@
Applying Lemma 2.6 for three times with largest possible multiplicities proves the claim. Indeed, each time the reductor $k=3d-\sum_{i=1}^{5}m_{i}$ affects the coefficients of $m$ in $m_{i}$ and $d$ significantly. In Table 1 we capture the coefficient of $m$ in $k$, and change of coefficients of $m$ in $m_{i}$ and $d$. The change in the constant term does not affect so we omit them in the table.
@@
Table 1. Reduction table for $I\left((20m)^{\times 9},(30m)^{\times 1}\right)_{36m-1}$
From the last row of Table 1, since, the coefficient of $m$ in the degree $d$ is less than that of the fifth point, we conclude that $I\left((20m)^{\times 9},(30m)^{\times 1}\right)_{36m-1}=0,\forall m$. Now, by applying the Lemma 2.7 one time, we get $I((20m)^{\times 15})_{36m-1}=0$, for all $m$, therefore, $\widehat{\alpha}(I)\geqslant\widehat{\alpha}(15)\geqslant\dfrac{36}{20}=\dfrac{5+4}{5}$.
@@
We claim that $I\left((20m)^{\times 1},(30m)^{\times 11}\right)_{48m-1}=0,\forall m.$ Similarly, we apply Lemma 2.6 for two times with largest possible multiplicities proves the claim. In Table 2 we capture the coefficient of $m$ in $k$, and change of coefficients of $m$ in $m_{i}$ and $d$. Again, the change in the constant term does not affect so we omit those in the table.
@@
Table 2. Reduction table for $I\left((20m)^{\times 1},(30m)^{\times 11}\right)_{48m-1}$
From the last row of Table 2, since, the coefficient of $m$ in the degree $d$ is less than that of the first point, we conclude that $I\left((20m)^{\times 1},(30m)^{\times 11}\right)_{48m-1}=0,\forall m.$ Now, by applying the Lemma 2.7 eleven times, we get $I((20m)^{\times 67})_{48m-1}=0$, for all $m$, therefore, $\widehat{\alpha}(I)\geqslant\widehat{\alpha}(67)\geqslant\dfrac{48}{20}=\dfrac{5+7}{5}$.
@@
Combined results in [BGHN22a, DSS24, NT19], Demailly’s Conjecture when $m=2$ now holds for any set of very general points, any set of at least $2^{N}$ general points, and any set of points in ${\mathbb{P}}^{3},{\mathbb{P}}^{4}$, and ${\mathbb{P}}^{5}$, except for $10\leqslant s\leqslant 13$ in ${\mathbb{P}}^{5}$. The case $10\leqslant s\leqslant 13$ in ${\mathbb{P}}^{5}$ is not known.